# DELIFT: Data Efficient Language model Instruction Fine-Tuning

### v2

@@ 4 Experimental Results
We conducted extensive experiments to evaluate the effectiveness of DELIFT across various fine-tuning scenarios, model scales, and datasets. This section details our experimental setup, baselines, evaluation metrics, and results analysis.
@@
Table 1: Results on Use Case 1: MixInstruct. Bold indicates the best performance. There is a 10.44% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 2.27% performance percentage drop from DELIFT to the next best baseline.
@@
Table 2: Results on Use Case 1: P3. Bold indicates the best performance. There is only a 0.76% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 3.23% performance percentage drop from DELIFT to the next best baseline.
4.1 Datasets and Use Cases
We evaluated DELIFT across the three previously described fine-tuning scenarios:
Use Case 1: Instruction Tuning We evaluated the effectiveness of DELIFT for use case 1 on two datasets: MixInstruct (Jiang et al., 2023) and P3 (Public Pool of Prompts) (Sanh et al., 2021). We randomly selected 21,000 train, 6,000 valid, and 3,000 test samples. Using the Facility Location (FL) submodular function, we aimed to select a subset of training data that was both representative and informative.
Use Case 2: Task-Specific Fine-Tuning We evaluated DELIFT for task-specific fine-tuning using two dataset pairs: (1) HotpotQA (Yang et al., 2018) with MMLU (Hendrycks et al., 2021), and (2) MixInstruct with MT-Bench (Zheng et al., 2023). We used the Facility Location Mutual Information (FLMI) submodular function to select the most informative samples from the training datasets (HotpotQA and MixInstruct) that shared relevant information with the target datasets (MMLU and MT-Bench, respectively).
Use Case 3: Continual Fine-Tuning We evaluated DELIFT in a continual fine-tuning setting using two dataset pairs: (1) SQuAD (Rajpurkar et al., 2016) paired with HotpotQA for general question-answering, and (2) proprietary query rewriting datasets covering IBM and government domains. [^1] In this task, non-standalone questions –questions that require previous context to answer– must be rewritten to be standalone. For example, ”How much is it?” should be rewritten to ”How much is the subscription for IBM Cloud?” Such queries are common in user-agent conversations where a user asks a follow-up to an agent. Our goal was to integrate new knowledge efficiently while minimizing redundancy. We employed the Facility Location Conditional Gain (FLCG) submodular function, selecting complementary samples from the new dataset (HotpotQA and Government query rewrite) that provided additional, non-overlapping information to the existing dataset (SQuAD and IBM query rewrite).
4.2 Experimental Setup
Models: We evaluated DELIFT on two state-of-the-art open-source models: Phi-3-mini-128k-instruct (Abdin et al., 2024): 3.8B parameters, Qwen2-72B-Instruct (Yang et al., 2024): 72B parameters. These models were chosen to demonstrate effectiveness across different model scales.
Model Qwen2 Phi-3 Method QLoRA QLoRA Initial 82.10 69.10 Random 79.31 65.16 SelectIT 79.13 65.24 LESS 80.35 66.72 DELIFT (SE) 80.10 66.36 DELIFT 81.70 68.70 Full Data 78.36 64.50
Table 3: Results on Use Case 2: HotpotQA and MMLU (5-shot) for Qwen2 and Phi-3 models (classification accuracy). Bold indicates the best performance. For Qwen2, DELIFT outperforms Full Data by 3.34%, while for Phi-3, it improves by 4.20%.
@@ 4.2 Experimental Setup
Table 4: Results on Use Case 2: MixInstruct and MT-Bench. Bold indicates the best performance. There is a 2.91% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 1.14% performance percentage drop from DELIFT to the next best baseline.
Metrics: We use a variety of metrics to characterize performance. For $n$-gram word overlap we use ROUGE (Lin, 2004). For semantic similarity we calculate the dot product between the embeddings from the bge-large-en-v1.5 model (Xiao et al., 2023); the embeddings are normalized to unit vectors, hence the closer the dot product is to 1, the more semantically similar the vectors (the metric is referred to as ‘BGE’). Additionally, we use Prometheus (Kim et al., 2023), specifically the prometheus-7b-v2.0 model, as an LLM-as-a-Judge (referred to as ‘LAJ’). With our custom rubric outlined in Appendix B, Prometheus assigns scores in a range of 1 to 5 (higher scores indicate better performance.) Finally, we use classification accuracy to evaluate MMLU.
Baselines: We evaluated DELIFT by comparing it against several baselines to understand its effectiveness in data selection. These baselines included: (1) SelectIT (Liu et al., 2024), which selects data using model feedback at the token, sentence, and model levels to identify useful samples; (2) LESS (Xia et al., 2024), which leverages LoRA approximated gradient-based influence estimation to prioritize impactful data points; (3) Random, which selects a fixed percentage (x%) of the dataset randomly, providing a benchmark for non-strategic selection; (4) DELIFT with Sentence Embedding Features (SE), which uses DELIFT but substitutes sentence embeddings as the feature space, employing a model-independent, pairwise similarity kernel instead of the utility kernel for submodular optimization; and (5) Full Data, where the entire dataset is used for fine-tuning, serving as an upper benchmark for performance. For In-Context Learning (ICL), the selected subsets from each baseline were used as the pool of examples, allowing us to evaluate how effectively each method supports ICL by providing relevant and informative data.
@@
Table 5: Results on Use Case 3: IBM and Government. Bold indicates the best performance. There is only a 0.31% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 3.89% performance percentage drop from DELIFT to the next best baseline.
@@
Table 6: Results on Use Case 3: SQuAD and HotpotQA. Bold indicates the best performance. There is only a 1.94% performance percentage drop from Full Data to DELIFT after pruning 70% of the data, and a 2.78% performance percentage drop from DELIFT to the next best baseline.
4.3 Results and Analysis
To ensure a fair and comprehensive evaluation of DELIFT, we conducted experiments across three distinct fine-tuning scenarios: instruction tuning, task-specific fine-tuning, and continual fine-tuning. For all subset selection methods—including DELIFT, Random, SelectIT, LESS, and DELIFT with Sentence Embdedding Features (SE)—we consistently selected 30% of the dataset as a subset, enabling direct comparisons between methods and with the full dataset baseline (see Section 4.4 for an ablation study examining the impact of subset size).
Use Case 1: Instruction Tuning Our first set of experiments focused on instruction tuning, a crucial task to enhancing a model’s ability to follow diverse instructions. As shown in Tables 1 and 2, DELIFT achieved a minimal performance drop of only 5.60% compared to using the full dataset while reducing the dataset by 70%. This demonstrates DELIFT’s capability to retain the most informative samples essential for instruction tuning. Furthermore, DELIFT outperformed other subset selection methods, achieving a 2.74% improvement and a substantial 26.21% advantage over the next best and worst-performing baselines, respectively. These results underscore DELIFT’s superior ability to maintain high performance with significantly reduced data, highlighting its efficacy in instruction tuning.
Use Case 2: Task-Specific Fine-Tuning In the task-specific fine-tuning scenario, we evaluated DELIFT using two dataset pairs: (1) HotpotQA (Yang et al., 2018) with MMLU (Hendrycks et al., 2021), and (2) MixInstruct paired with MT-Bench (Zheng et al., 2023). Results, presented in Tables 4 and 3, demonstrate DELIFT’s consistent and competitive performance across different task pairs. A particularly noteworthy outcome emerged from the HotpotQA-MMLU pair, where DELIFT not only matched but exceeded the performance of the full dataset, achieving a 5.51% improvement. This indicates that DELIFT’s selective approach can effectively filter out noise and focus on the most relevant and informative samples, yielding enhanced task-specific adaptation even with reduced data.
Use Case 3: Continual Fine-Tuning The third use case examined DELIFT’s efficacy in continual fine-tuning, where models need to incorporate new information while retaining previously learned knowledge. As detailed in Tables 5 and 6, DELIFT demonstrated remarkable consistency, showing only a marginal 1.13% performance drop compared to using the full dataset. Moreover, DELIFT outperformed the second-best baseline by 3.33% and the worst baseline by 23.88%, highlighting its superiority in data selection. In specialized tasks such as query rewriting, DELIFT even surpassed the performance of the full dataset, suggesting that its selective approach effectively prunes noisy or irrelevant data points, thereby enhancing model performance.
4.4 Ablation Study: Impact of Subset Size
To assess how subset size influences DELIFT’s performance, we conducted an ablation study by varying the subset size from 5% to 100% of the full dataset across three use cases. The results, detailed in Appendix A and illustrated in Figure 2, show that LAJ scores generally increase with subset size. Utilizing the full dataset consistently yields the highest performance, highlighting the benefits of larger training sets. However, for methods such as DELIFT, SelectIT, and LESS, performance gains plateau or slow beyond a 50% subset size, indicating that additional data offers minimal benefits and may introduce redundancy. Importantly, DELIFT outperforms all baselines across subset sizes from 5% to 100%, demonstrating its robustness and effectiveness in selecting informative samples regardless of subset size. These findings suggest that carefully selected smaller datasets can achieve comparable performance to larger, unfiltered datasets, which is particularly valuable for resource-intensive large language models.
4.5 Discussion
The comprehensive results across all three use cases highlight DELIFT’s effectiveness and versatility. By consistently reducing data requirements by up to 70% while maintaining—and in some cases improving—performance, DELIFT addresses a critical challenge in large language model fine-tuning. The superior performance of DELIFT can be attributed to its novel pairwise utility metric and the use of tailored submodular functions for each fine-tuning stage. This approach enables DELIFT to select not only representative and diverse samples but also to reduce noise present in the full dataset. The ability to outperform full datasets in certain scenarios, particularly in niche tasks like query rewriting, underscores DELIFT’s capacity to distill the most relevant and informative data points. These findings have significant implications for the accessibility and efficiency of LLM fine-tuning. By dramatically reducing the amount of data required for effective fine-tuning, DELIFT paves the way for more widespread adoption and application of large language models across various domains, especially in resource-constrained environments. Furthermore, DELIFT’s consistent outperformance of existing data selection techniques across various fine-tuning scenarios and model scales demonstrates its robustness and broad applicability, making it a valuable tool for researchers and practitioners alike. In conclusion, our experimental results firmly establish DELIFT as a powerful and efficient method for data selection in LLM fine-tuning. By addressing the critical challenge of optimal data selection, DELIFT not only enhances the efficiency of model training but also opens new possibilities for fine-tuning large language models in domains where data or computational resources may be limited.
@@ 6 Acknowledgement
This work used the Delta system at the National Center for Supercomputing Applications through allocation CIS240550 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
Appendix A Subset Size Comparison

### v3

@@ 4 Experimental Results
We conduct extensive experiments to evaluate DELIFT in three fine-tuning scenarios: (1) Instruction Tuning, (2) Task-Specific Fine-Tuning, (3) Continual Fine-Tuning. This section outlines our experimental setup and baselines, followed by results across multiple models, datasets, and fine-tuning paradigms, along with interpretative commentary to utilize the extra space for deeper insights.
4.1 Experimental Setup
Models.
We evaluate DELIFT on both base LLMs (Llama-3.2-3B, Mistral-7B-v0.1, opt-125m) and instruction-tuned LLMs (Qwen2-72B-Instruct, Phi-3-mini-128k-instruct). This variety covers different parameter scales (millions vs. billions) along with diverse adaptation strategies (ICL, QLoRA, and, for smaller models, full fine-tuning). This diversity in model size and training setup provides a robust test-bed for assessing DELIFT across both resource-rich and more constrained environments.
Datasets.
We group the datasets by the primary goal of fine-tuning, ensuring a clear mapping from the data to the corresponding submodular objective. In particular: 1. Instruction Tuning: Mix-Instruct (Jiang et al., 2023), P3 (Sanh et al., 2021). Both aim to enhance general instruction-following behavior, featuring a variety of task prompts and user requests. 2. Task-Specific Fine-Tuning: HotpotQA (Yang et al., 2018) aligned with MMLU (Hendrycks et al., 2021), Mix-Instruct aligned with MT-Bench (Zheng et al., 2023), and Mix-Instruct aligned with GSM-8k (Cobbe et al., 2021). These pairings allow us to extract only the most relevant samples from a large corpus to improve performance on a specific target benchmark. 3. Continual Fine-Tuning: (a) SQuAD (Rajpurkar et al., 2016) paired with HotpotQA to inject more complex, multi-hop reasoning data after simpler QA, and (b) a proprietary IBM/Government domain query rewriting dataset. [^1] Query rewriting transforms follow-up queries (e.g., “How much is it?”) into standalone forms (e.g., “How much is the subscription for IBM Cloud?”). This is essential for real-world systems where user queries often rely on prior context.
In all cases, we fixed an approximate budget of 30% for subset selection unless otherwise noted, striking a balance between data efficiency and coverage.
Baselines.
In addition to a Full Data baseline, where the entire training set is used, we compare DELIFT with:
1. Random: Uniformly selects 30% of the training set to provide a simple, model-agnostic benchmark.
2. SelectIT (Liu et al., 2024): Generates self-reflection prompts within the LLM to rate data quality, filtering out lower-quality samples.
3. LESS (Xia et al., 2024): Employs gradient-based influence estimates, approximated via LoRA, to identify highly impactful data points for model parameter updates.
4. DEFT-UCS: Uses sentence embeddings to cluster the dataset for diversity; although it captures semantic variety, it lacks explicit model feedback to guide selection.
4. DELIFT (SE): Operates the same submodular optimization as DELIFT but replaces our utility-based kernel with static sentence embedding similarities to highlight the benefit of DELIFT’s dynamic, model-aware approach.
These baselines range from naive (Random) to sophisticated (gradient-based, reflection-based, or embedding-based), furnishing a thorough comparison against DELIFT.
Metrics.
We measure model performance across different aspects:
1. ROUGE (Lin, 2004): Focuses on $n$-gram overlap for summarization tasks or generative text alignment.
2. BGE (Xiao et al., 2023): Evaluates semantic similarity through the dot product of normalized sentence embeddings.
3. LAJ (LLM-as-Judge) (Kim et al., 2023): Assigns a 1–5 rating reflecting correctness, clarity, and instruction adherence; we detail the scoring rubric and example prompts in Appendix C.
4. Classification Accuracy: Used primarily for multiple-choice tasks like MMLU.
By combining these metrics, we obtain both text-overlap measures (ROUGE), semantic evaluations (BGE), a holistic LLM-based score (LAJ), and a classification perspective (accuracy), offering a comprehensive view of model improvements under each fine-tuning scenario.
4.2 Use Case 1: Instruction Tuning
Setting.
Instruction tuning aims to broaden a model’s capacity to follow instructions spanning diverse domains and styles. We adopt the Facility Location (FL) objective to maximize coverage of varied instruction types.
4.2.1 Results on Instruction-Tuned Models.
Tables 1 and 2 compare DELIFT with baselines on Mix-Instruct and P3, using two instruction-tuned models (Qwen2-72B-Instruct and Phi-3-mini-128k-instruct). Across both ICL and QLoRA, DELIFT surpasses other subset selection strategies and manages to prune up to 70% of data with minimal loss. Interestingly, DELIFT approaches—or occasionally exceeds—the Full Data baseline, indicating strong redundancy in instruction data.
@@ 4.2 Use Case 1: Instruction Tuning
Table 1: Use Case 1: MixInstruct. Bold indicates best performance. DELIFT prunes 70% data yet stays close to Full Data.
@@
Table 2: Use Case 1: P3. DELIFT again discards most data while retaining strong performance.
4.2.2 Evaluation on a Base Model.
Table 3 verifies these findings on a non-instruction-tuned Llama-3.2-3B (base). Even without prior instruction alignment, DELIFT outperforms gradient- and embedding-based baselines by a solid margin, corroborating that the utility metric is robust to model initialization.
Method ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ Initial 25.88 61.97 1.90 28.51 73.14 2.48 Random 28.64 77.14 2.59 38.79 79.89 2.53 SelectIT 42.67 81.54 2.67 45.87 84.67 2.61 DEFT-UCS 41.55 80.12 2.63 41.03 81.86 2.59 LESS 44.99 82.48 2.69 50.54 84.14 2.78 DELIFT (SE) 51.19 83.54 2.72 55.32 85.92 2.79 DELIFT 54.58 88.29 2.83 58.57 90.98 2.98 Full Data 55.46 92.71 3.31 61.23 95.52 3.10
Table 3: Use Case 1: Llama-3.2-3B (base) on Mix-Instruct (30% subset).
Takeaway (Use Case 1): The FL objective and utility-based kernel consistently prune large volumes of instruction data without undermining final performance—a strong indicator that many instruction examples are either repetitive or uninformative.
4.3 Use Case 2: Task-Specific Fine-Tuning
Setting.
We refine models for specialized domains (reasoning, QA). Using Facility Location Mutual Information (FLMI), we select examples aligned with a target dataset $\mathcal{D}_{T}$.
4.3.1 HotpotQA $\to$ MMLU & MixInstruct $\to$ MT-Bench.
Method Qwen2 (QLoRA) Phi-3 (QLoRA) Initial 82.10 69.10 Random 79.31 65.16 SelectIT 79.13 65.24 LESS 80.35 66.72 DELIFT (SE) 80.10 66.36 DELIFT 81.70 68.70 Full Data 78.36 64.50
Table 4: Use Case 2: HotpotQA and MMLU (5-shot) for Qwen2 and Phi-3 (classification accuracy). DELIFT outperforms Full Data by 3.34% for Qwen2 and by 4.20% for Phi-3.
@@ 4.3 Use Case 2: Task-Specific Fine-Tuning
Table 5: Use Case 2: Mix-Instruct and MT-Bench. Bold indicates the best performance. There is a 2.91% drop from Full Data to DELIFT after pruning 70% of the data, and a 1.14% drop from DELIFT to the next best baseline.
Empirical Findings.
Tables 5 and 4 illustrate DELIFT’s consistent success in bridging the gap between diverse source datasets (Mix-Instruct or HotpotQA) and target benchmarks (MT-Bench, MMLU). In some cases (HotpotQA $\to$ MMLU), DELIFT surpasses the Full Data baseline by pruning unhelpful training points.
4.3.2 Further Experiments on Complex Reasoning Tasks.
To rigorously test DELIFT on more challenging datasets, we apply it to Mistral-7B-v0.1 (base) using Mix-Instruct aligned with GSM-8k, a corpus of math word problems that demand multi-step logical reasoning. Table 6 shows that DELIFT yields strong improvements under both ICL and QLoRA, surpassing all competing baselines. Even the simpler FL-only variant outperforms LESS and DEFT-UCS, highlighting the advantage of carefully selected examples when dealing with more complex reasoning tasks. Moreover, using FLMI for task alignment confers an additional boost, reinforcing the importance of matching the submodular objective to the dataset’s complexity and problem-solving requirements.
Method ICL QLoRA ROUGE BGE LAJ ROUGE BGE LAJ Initial 17.06 55.13 1.35 27.95 65.92 1.36 Random 19.18 56.03 1.38 33.85 81.74 1.85 SelectIT 30.27 77.23 1.45 42.29 88.17 2.49 DEFT-UCS 30.06 76.99 1.59 41.45 87.84 2.08 LESS 31.69 77.87 2.26 43.86 88.22 2.60 DELIFT (SE) 31.84 78.62 2.44 43.04 90.53 2.54 DELIFT (FL instead of FLMI) 32.30 78.54 2.54 44.62 91.04 2.63 DELIFT 33.25 79.10 2.56 46.12 92.97 2.71 Full Data 35.33 81.59 2.79 49.10 94.57 2.85
Table 6: Use Case 2: Mistral-7B-v0.1 (base): Mix-Instruct $\to$ GSM-8k with FLMI.
Takeaway (Use Case 2): By explicitly matching data to a target domain or benchmark, DELIFT discards irrelevant samples and—in some cases—yields better results than training on the entire dataset.
4.4 Use Case 3: Continual Fine-Tuning
Setting.
A model must assimilate new data without forgetting old knowledge. We use Facility Location Conditional Gain (FLCG) to prioritize novel samples while avoiding overlaps.
4.4.1 Examples: IBM $\to$ Government, SQuAD $\to$ HotpotQA.
Tables 7 and 8 confirm that DELIFT mitigates redundancy. Notably, on IBM $\to$ Government, DELIFT sees less than 1% drop vs. Full Data, versus a 3–4% gap for baselines like LESS and DEFT-UCS.
@@ 4.4 Use Case 3: Continual Fine-Tuning
Table 7: Use Case 3: IBM and Government. DELIFT effectively selects novel examples, outperforming baselines.
@@
Table 8: Use Case 3: SQuAD and HotpotQA. Bold indicates best performance. There is only a 1.94% drop from Full Data to DELIFT after pruning 70% of the data, and a 2.78% drop from DELIFT to the next best baseline.
Takeaway (Use Case 3): By prioritizing complementary rather than redundant data, DELIFT preserves previous capabilities and focuses on new insights.
4.5 Further Comparisons and Ablations
4.5.1 Full Fine-Tuning vs. QLoRA.
We also validated DELIFT under full fine-tuning on a smaller opt-125m (Table 9). Results show consistent gains for DELIFT, indicating that the utility-based selection approach is independent of the particular fine-tuning methodology.
Method QLoRA Full Fine-Tuning ROUGE BGE LAJ ROUGE BGE LAJ Initial 9.04 40.50 1.19 9.57 40.86 1.14 Random 12.55 46.99 1.25 13.07 47.91 1.30 SelectIT 14.78 49.80 1.26 15.35 50.42 1.28 DEFT-UCS 15.12 50.29 1.39 15.16 50.29 1.39 LESS 15.52 50.70 1.38 16.02 51.30 1.40 DELIFT (SE) 18.81 55.02 1.35 17.06 55.93 1.38 DELIFT 19.72 57.98 1.43 19.87 58.11 1.45 Full Data 20.39 60.39 1.95 21.64 61.70 2.05
Table 9: opt-125m on Mix-Instruct, comparing QLoRA vs. full fine-tuning.
4.5.2 Comparing Submodular Objectives.
Although FL alone can beat naive baselines in specialized or incremental settings (see Table 6), the specialized objectives (FLMI for domain tasks, FLCG for continual updates) yield stronger alignment. This underscores the importance of matching the submodular objective to the fine-tuning stage.
4.5.3 Ablation on Subset Size.
Beyond consistently using 30% of the data in our main experiments, we investigated how varying the subset size influences performance. We tested budgets ranging from as little as 5% up to 50% of the original training set (in increments of 10%). As shown in Figure 2, performance under all methods generally improves with larger subsets but exhibits diminishing returns beyond 30–40%. Notably, DELIFT consistently yields higher LAJ scores than other baselines at every subset size, demonstrating its robustness even under very aggressive pruning (e.g., 5%). These findings reinforce that substantial reductions in training data are possible without significant performance degradation, and that DELIFT identifies highly informative samples more effectively than competing methods.
4.6 Key Observations
Our experiments reveal four consistent patterns across datasets, model scales, and adaptation methods:
  1. 1.
    Utility-based kernel outperforms static or gradient-based methods, indicating that measuring how one sample improves predictions on another is a stronger signal than mere embeddings or approximate gradients.
  2. 2.
    Stage-specific objectives (FL, FLMI, FLCG) are crucial for addressing the unique needs of instruction tuning (diverse coverage), task-specific alignment, and continual complementarity.
  3. 3.
    Significant pruning (up to 70%) is routinely possible, often with minimal or zero performance deterioration, suggesting vast redundancy in large corpora.
  4. 4.
    Method-agnostic gains, as DELIFT consistently improves results under ICL, QLoRA, and even full fine-tuning on smaller models.
In the next section, we summarize these findings, discuss limitations such as the $O(n^{2})$ cost of utility computation, and propose directions for future work, including adaptive selection during training, extensions to multimodal data, and fairness considerations.
@@ 6 Acknowledgement
A part of this work used the Delta system at the National Center for Supercomputing Applications through allocation CIS240550 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
Appendix A Theoretical Foundations and Connections Between the Utility Metric and Information Theory
Theorem 2
Let $y_{i}=(y_{i1},y_{i2},\dots,y_{iT})$ be a sequence of tokens with ground truth distribution $GT_{i}$, where $GT_{i}$ assigns probability 1 to the sequence $y_{i}$ and 0 to all other sequences. Let $p(y_{i}\mid x_{i})$ be the predicted probability of $y_{i}$ given input $x_{i}$, and $p(y_{i}\mid x_{i},x_{j},y_{j})$ be the predicted probability of $y_{i}$ given $x_{i}$ and an in-context example $(x_{j},y_{j})$. Define the utility metric $UF_{ij}$ using a general distance metric $d(\cdot,\cdot)$ between probability distributions:
$UF_{ij}=d(GT_{i},\,p(y_{i}\mid x_{i}))-d(GT_{i},\,p(y_{i}\mid x_{i},x_{j},y_{j})).$
Claim: When the distance metric $d(\cdot,\cdot)$ is the Kullback-Leibler divergence $D_{\text{KL}}$, the utility metric $UF_{ij}$ is equal to the pointwise mutual information (PMI) between the sequence $y_{i}$ and the in-context example $(x_{j},y_{j})$ conditioned on $x_{i}$:
$UF_{ij}=\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i})=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}.$
Furthermore, $UF_{ij}$ can be expressed as the sum of conditional PMI over the tokens in $y_{i}$:
$UF_{ij}=\sum_{t=1}^{T}\operatorname{PMI}\big{(}y_{it};\,x_{j},y_{j}\mid x_{i},y_{i,<t}\big{)},$
where $y_{i,<t}=(y_{i1},y_{i2},\dots,y_{i(t-1)})$ denotes the sequence of previous tokens up to position $t-1$.
Proof:
1. Computing KL-Divergence Between Ground Truth and Predicted Distributions:
Since $GT_{i}$ assigns probability 1 to the specific sequence $y_{i}$, the KL-divergence simplifies as follows:
$d(GT_{i},\,p(y_{i}\mid\cdot))=D_{\text{KL}}(GT_{i}\parallel p(y_{i}\mid\cdot))=-\log p(y_{i}\mid\cdot),$
because the KL-divergence between a one-hot distribution and any other distribution reduces to the negative log-probability of the assigned event.
2. Computing the Utility Metric $UF_{ij}$:
The utility metric becomes:
$UF_{ij}$ $=-\log p(y_{i}\mid x_{i})+\log p(y_{i}\mid x_{i},x_{j},y_{j})$
$=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}.$
3. Expressing $UF_{ij}$ as Pointwise Mutual Information:
The conditional pointwise mutual information between $y_{i}$ and $(x_{j},y_{j})$ given $x_{i}$ is defined as:
$\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i})=\log\frac{p(y_{i},x_{j},y_{j}\mid x_{i})}{p(y_{i}\mid x_{i})\,p(x_{j},y_{j}\mid x_{i})}.$
Using the chain rule:
$p(y_{i},x_{j},y_{j}\mid x_{i})=p(y_{i}\mid x_{i},x_{j},y_{j})\,p(x_{j},y_{j}\mid x_{i}).$
Substituting back:
$\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i})$ $=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})\,p(x_{j},y_{j}\mid x_{i})}{p(y_{i}\mid x_{i})\,p(x_{j},y_{j}\mid x_{i})}$
$=\log\frac{p(y_{i}\mid x_{i},x_{j},y_{j})}{p(y_{i}\mid x_{i})}.$
Therefore:
$UF_{ij}=\operatorname{PMI}(y_{i};\,x_{j},y_{j}\mid x_{i}).$
4. Expanding $UF_{ij}$ as Sum of Conditional PMI Terms:
We expand $p(y_{i}\mid\cdot)$ using the chain rule:
$p(y_{i}\mid\cdot)=\prod_{t=1}^{T}p(y_{it}\mid\cdot,y_{i,<t}),$
where $y_{i,<t}$ is the sequence of previous tokens up to time $t-1$.
Substituting back into $UF_{ij}$:
$UF_{ij}$ $=\log\frac{\prod_{t=1}^{T}p(y_{it}\mid x_{i},x_{j},y_{j},y_{i,<t})}{\prod_{t=1}^{T}p(y_{it}\mid x_{i},y_{i,<t})}$
$=\sum_{t=1}^{T}\left[\log p(y_{it}\mid x_{i},x_{j},y_{j},y_{i,<t})-\log p(y_{it}\mid x_{i},y_{i,<t})\right]$
$=\sum_{t=1}^{T}\operatorname{PMI}\big{(}y_{it};\,x_{j},y_{j}\mid x_{i},y_{i,<t}\big{)}.$
This shows that $UF_{ij}$ is the sum of the conditional PMI of each token $y_{it}$ with $(x_{j},y_{j})$ given $x_{i}$ and the previous tokens.
Conclusion:
When $d(\cdot,\cdot)=D_{\text{KL}}$, the utility metric $UF_{ij}$ precisely equals the conditional PMI between $y_{i}$ and $(x_{j},y_{j})$ given $x_{i}$.
A.1 Why Euclidean Distance is Preferred Over KL-Divergence for Subset Selection
The effectiveness of subset selection algorithms, including facility location functions, depends critically on the properties of the chosen distance metric d(·,·). Euclidean distance offers several key advantages over KL-divergence for this purpose:
  1. 1.
    Mathematical Properties
    • Euclidean distance is non-negative, finite, and symmetric (d(a,b) = d(b,a))
    • KL-divergence can be infinite or undefined with zero probabilities and lacks symmetry $D_{KL}(P\parallel Q)\neq D_{KL}(Q\parallel P)$
  2. 2.
    Computational Advantages
    • Euclidean distance uses simple arithmetic operations (subtraction, squares, square roots)
    • KL-divergence requires more complex logarithmic calculations and division operations
  3. 3.
    Robustness in Practice
    • Euclidean distance handles zero probabilities gracefully
    • KL-divergence becomes undefined with zero probabilities, which occur frequently in real data
Impact on Subset Selection: The facility location function requires positive, finite similarity measures to model coverage effects accurately. Euclidean distance satisfies these requirements, while KL-divergence’s potential negative or infinite values can disrupt optimization.
Conclusion: While KL-divergence offers theoretical connections to mutual information, Euclidean distance provides:
  • Guaranteed positive and finite utility metrics
  • Superior computational efficiency
  • Better numerical stability
These practical advantages make Euclidean distance the preferred choice for computing the utility metric $UF_{ij}$ in subset selection algorithms.
Appendix B Subset Size Comparison